Trying to solve Y=-2X^2+BX+5 Assume max value of X=2 Find B
I'm having problem expressing the problem in the Y=A(X-H)^2+K form. The first step of getting it to Y=-2(X-2)^2+K is where I'm having trouble. I can solve the problem once it is expressed in the form. I just don't know how to change one form to the other. I'm trying to brush up so I can help my daughter on her SAT practice tests. Thank you very much for any assistance you might provide.

Answer by nerdybill(7384)

Trying to solve Y=-2X^2+BX+5 Assume max value of X=2 Find B
.
If the "max value" is achieved when x=2 then
x=2 is also the "axis of symmetry"
.
"axis of symmetry" formula:
x = -b/(2a)
(THIS, you have to remember--it is derived from the quadratic formula)
plugging in the given values:
2 = -b/(2*(-2))
2 = -b/(-4)
2 = b/4
8 = b